Two products

YR 1

YR 2

P1

W1

P2

W2

Call center

10

50

12

50

Banking

10

50

12

60

d. Using year 1 prices, what is real GDP per worker in year 1 and year 2? What is labour productivity growth between year 1 and year 2 for the whole economy? Now suppose that banking services in year 2 are not the same as banking services in year 1 because they include internet banking, which year 1 banking services didn’t include. The technology for internet banking was available in year 1 but the price of banking services with internet banking in year 1 was $13 and no one chose that package. However, in year 2 the price of banking services with internet banking was $12 and everyone chose to have that package in year 2 (that is, in year 2 no one chose to have the year 1 banking services package without internet banking). 


1
Expert's answer
2021-12-03T11:50:03-0500

(a)

From the table:

Nominal GDP

Year 1

call center services:

"=10\times100=1000"

Banking services:

"=10\times200=2000"

Total nominal GDP:

"=3000"

Year 2

call center services:

"=12\times100=1200"

Banking services:

"=12\times230=2760"

Total nominal GDP:

"=3960"

(b)

Real GDP per worker

year 1:

"=\frac{3000}{100}=30"

Year 2:

"=\frac{3960}{110}=36"

(c)

Real GDP for year 2

Call center services:

"=10\times100=1000"

Banking services:

"=13\times 230=2990"

Total Real GDP for year 2:

"=3990"

(d)

"GDP deflator=\frac{Nominal GDP}{Real GDP}\times 100"

"=\frac{3960}{3990}\times100=99.25"